Circular No. 6507

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

4U 1820–30 A. P. Smale, Goddard Space Flight Center (GSFC), NASA and Univer-

sities Space Research Association; and W. Zhang and N. E. White, GSFC, report: “RXTE observations of the low-mass x-ray binary 4U 1820–30 performed on Oct. 15 and Oct. 26–30 reveal the presence of high-frequency quasiperiodic oscillations (QPOs) when the source is in its low state (160– 200 mCrab). The centroid frequency of the main peak varies between 546 ± 2 and 796 ± 6 Hz, and is tightly correlated with source intensity. The QPO width is typically 20–30 Hz, with rms amplitude 5.0 ± 0.5 percent. When the source is at the upper end of this range, a second significant QPO peak appears with frequency 1066 ± 5 Hz, width 46 ± 12 Hz, and rms amplitude 3.2 ± 0.5 percent. When the source brightens above 200 mCrab, neither QPO is detected.”

COMET C/1995 O1 (HALE-BOPP) Extension to the ephemeris on IAUC 6440, from orbital elements by

B. G. Marsden on MPC 28052:

1996/97 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Nov. 23 17^{h}56.^{m}41 − 1^{o}30.^{′}5 2.981 2.245 35.^{o}0 14.^{o}6 3.9
28 18 01.13 − 0 58.1 2.948 2.182 32.6 14.1 3.7
Dec. 3 18 06.18 − 0 22.0 2.909 2.120 30.5 13.7 3.6
8 18 11.58 + 0 18.3 2.864 2.057 28.9 13.4 3.4
13 18 17.31 + 1 03.1 2.812 1.994 27.7 13.3 3.2
18 18 23.40 + 1 53.2 2.755 1.930 27.0 13.4 3.1
23 18 29.86 + 2 49.0 2.691 1.867 26.9 13.8 2.9
28 18 36.71 + 3 51.3 2.621 1.804 27.2 14.4 2.7
Jan. 2 18 43.99 + 5 00.8 2.546 1.740 28.0 15.4 2.4
7 18 51.74 + 6 18.4 2.466 1.677 29.2 16.6 2.2
12 19 00.02 + 7 45.2 2.381 1.614 30.6 18.1 2.0
17 19 08.92 + 9 22.4 2.291 1.551 32.3 19.8 1.7
22 19 18.54 +11 11.0 2.198 1.489 34.1 21.7 1.4
27 19 29.01 +13 12.7 2.102 1.428 36.0 23.9 1.2
Feb. 1 19 40.54 +15 28.7 2.004 1.368 37.9 26.3 0.9
6 19 53.38 +18 00.8 1.906 1.309 39.7 28.8 0.6
11 20 07.85 +20 50.1 1.807 1.251 41.4 31.5 0.3
1996 November 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green